Given three random variables Xi, X2, and X such that X[Xi X2 X 20 -1 3 1 0.5 1 E [X]-μ | 0 | and var(X)=Σー| 0 0.5 | com pute: 2 c) var(X2-X3 (d) var(X2 + X3) (e) cov(4X2 +X1,3Xi - 2X3)
Answer #1

TOPIC: Use of mean vector and the dispersion matrix of a random vector to find the values required.

Considem the mandom vecton es with, E (X) = 쓰 = 0 3 .5 2 we need, E (x, X) b) We need, E (3% -x2+4x3) = 3x1-0 +4x(-:). = 3-8· 1-51 e we need, var. (x2-Xs). = 3+1-2x0.53+1+2x0.5. 国 ene G) we have Conn (χ3tx2 , 2x2-35). Cov (ス3tX1, 2左-%)de have, var (x) : where, a constant ,l- 12 +18ニ30 Cov ov33) 2- Cav (x).xs)一So, 0 ) 10 (0 150 Rounded to wt decimal 5 x 30 places)